The network switches still need to be connected to the network, in the meantime, everything will be completed so they can be imaged
Thu, Jul 2
Rack A5 Port 30
@fgiunchedi icinga1001 is in rack C8, that is now a 10G rack. Do you still want this server there or can we move to another rack that is 1G only and eventually migrate icinga1001 to the same rack?
Tue, Jun 30
new-msw1-eqiad has the correct JUNOS 18.1.3 and the configuration has been copied. Currently connected to port 2 on the a8-scs and can be moved to the correct port once we 100% migrate to the new switch.
@Gehel I am getting a partman error. Is the partman recipe given correct? The raid10-4dev is not working.
Thu, Jun 25
@Dzahn Could you try to image one of these and let me know if you see a setting missed. I am able to login to the ilo via mgmt network but when I use the install script on cumin I am getting an IPMI connection error.
This is being worked on, I had to put the OS image back on the usb stick. When I reset the switch to factory default the usb was wiped as welll.
Mon, Jun 22
I am having an issue with these. While running the script it gives me an error about IPMI. These are HP servers and I do not know of any manual IPMI setting.
Fri, Jun 12
@wiki_willy I completely forgot but restbases have ssds that were purchased separately from the servers. I believe this is the task of the original purchase. T158795. Also, this is a past task for ordering ssds for Restbase-dev1006 https://phabricator.wikimedia.org/T224260.
removed from networks switch, all dns entries removed, removed from rack and netbox has been updated to offline. Resolving this task
Wed, Jun 10
@akosiaris Have any the initial steps been completed with this decom task?
All of these servers have been removed from the racks, networks switch and moved to offline in netbox
@fgiunchedi These are moved and installed. Resolving this task.
Tue, Jun 9
These are all finsihed
Mon, Jun 8
thanos-fe1003 is the only one installed at the moment.
Jun 4 2020
removed from rack, netbox updated, switch updated, removed dns (asset tag)
Jun 2 2020
the server is out of warranty, I reseated both CPUs and cleared the system event log. The server booted okay. I will resolve this for now, please open again if the issue comes back.
There is a larger issue with this server, replaced the CPU but noticed the power supplies are both failed. I could also smell burning in the server, swapped the power supplies with decom spares and they psu's started smoking and something definitely burned inside the power supply. This will be down for awhile
Jun 1 2020
This was out there lingering without the property tags. I found it and will finish the decom process.
network switches are updated but they are disabled until I set up bios and image
the AHS log has been uploaded to HP per their request. Looks like we're going to be okay with the warranty thing
Added to network switches and put in disabled vlan until bios is set up and ready for imaging.
Network switches have been updated but I disabled the ports until after the bios is set up and ready for imaging.
May 29 2020
removed from rack, dns removed, netbox updated.
Removed from rack, mgmt dns removed, netbox updated
Removed from rack, dns removed, removed from network switch cfg, updated netbox.
@wiki_willy I submitted a ticket with HPE, we'll see what they say
May 24 2020
I submitted a ticket with Dell for a replacement CPU. SR1025619583
May 21 2020
this has not been completed
Removed from rack, netbox updated
removed from rack, mgmt dns removed, switch ports were already removed. Updated netbox
May 18 2020
I removed from rack and updated netbox
I see on the procurement task that these were supposed to be here on the 6 May but I have not seen them. @Jclark-ctr have you received them?
updated and removed
Removed from rack, updated netbox, removed dns, verified no entries in site.pp or puppet.
May 13 2020
updated netbox, removed from rack.
verified all entries are gone, removed from rack, updated netbox to offline
Removed server from rack, updated netbox
Checked everything and all entries removed, server is off rack, removed the storage array attached to it and set to offline and piled for sale.
all dns and network switch records have been removed, servers have already been sold